OpenStudy (anonymous):

Solve for n: 6(5n – 1) = 2(15n – 3) -1 3 all real solutions no solution

OpenStudy (anonymous):

6 (5 n-1) = 2 (15 n-3)

OpenStudy (anonymous):

30n - 6 = 30n - 6 30n - 6 + 6 - 30 n = 0 So No solution

OpenStudy (anonymous):

yep no solution by solving it